# Masked Diffusion Models as Energy Minimization

### v1

@@ 1 Introduction
Masked diffusion models (MDMs) [4, 8, 34, 44, 36, 38] have emerged as a powerful class of generative models for discrete data. By reversing a stochastic masking process, MDMs iteratively generate sequences through a series of unmasking steps, guided by learned denoising functions. This simple yet flexible architecture has shown promising empirical performance across text generation [43], protein generation [39, 40], and image generation [15, 23, 47].
Despite their empirical success, the underlying principles that govern the sampling efficiency of MDMs—particularly in few-step regimes—remain poorly understood. Existing works typically adopt manually designed mask schedules (e.g., linear or sine) without theoretical justification. In contrast, in continuous domains [27, 22, 25] and discrete flow matching [37], recent work has drawn deep connections between diffusion models and optimal transport, motivating questions such as: Can MDMs be similarly understood through a similar lens? Can we characterize optimal sampling schedules in a principled way? How do these schedules relate to the geometry of the underlying probability space?
@@ 2.1 Discrete Probability Flows and Masked Diffusion Models
Discrete Probability Flows (DPFs). DPFs [16, 14, 32] define continuous-time transformations over structured distributions in finite state spaces. In text generation tasks, a state $z=(z^{1},...,z^{n})\in\mathcal{D}^{n}$ typically denotes a token sequence of fixed length $n$ drawn from a vocabulary $\mathcal{D}$ of size $d$. Formally, DPFs introduce a family of parameterized distributions $(p_{t}(z))_{t\in[0,1]}$ that interpolate between a tractable base distribution $p_{0}(z)$ (e.g., a uniform distribution) and a target distribution $p_{1}(z)$.
A DPF is governed by a time-dependent transition rate matrix $(Q_{t})_{t\in[0,1]}$, [^1] We adopt the MDM terminology rather than the “velocity functions” used in the flow matching literature. which specifies transition probabilities between states. However, the reverse implication does not hold: different rate matrices can induce the same DPF [32]. We will omit the range of the time index $t\in[0,1]$ for brevity.
Masked Diffusion Models (MDMs). MDMs [4, 8] represent a subclass of DPFs built upon absorbing Markov chains. These models capture two temporal processes: a masking process (operating backward in time, $t=1\rightarrow 0$) and its reverse, the unmasking process ($t=0\rightarrow 1$) as follows.
@@
$p_{t|1}(x^{i}|x_{1})=\alpha_{t}\delta_{[\mathrm{M}]}(x^{i})+(1-\alpha_{t})\delta_{x_{1}^{i}}(x^{i}),$ (3)
@@
as shown in existing work [44, 38]. Existing methods model the likelihood ratio $\frac{p_{t}(x)}{p_{t}(z)}$ in Eq. (2)—also known as the concrete score [19, 26] or its equivalent formulations [38, 36, 44]—and optimize evidence lower bounds (ELBO) of the log-likelihood [4]. In particular, the ELBO is invariant to the choice of $\alpha_{t}$ [12, 38, 36] by change of variables. We provide proofs for Eq. (4) in Appendix C.1 and the invariance of ELBO in Appendix C.2 for completeness.
@@ 2.2 Kinetic and Conditional Kinetic Energy
Kinetic energy [2, 13, 27, 37] provides a principled framework for quantifying the transport cost of probability flows, forming an optimization objective that yield improved sampling trajectories.
@@
The function $\gamma_{t}$ is a smooth, strictly increasing schedule satisfying the boundary conditions $\gamma_{0}=0$ and $\gamma_{1}=1$. Different choices of $\gamma_{t}$ implement various temporal weighting schemes for the DPF. This generalizes the unweighted kinetic energy [37], where $\gamma_{t}=t$, and facilitates our theoretical analysis.
The quadratic dependence on the transition rates $Q_{t}(z,x)$ in Eq. (5) mirrors the classical velocity-squared form of kinetic energy. As $Q_{t}$ controls sampling behavior, minimizing $\mathcal{E}_{k}$ seeks minimal-cost sampling paths between distributions [24, 22, 25], a central goal in efficient generative modeling and few-step sampling [2, 13, 27].
@@
Definition 2.2 extends Definition 2.1 to the conditional setting, in line with analogous energy formulations in the continuous domain [27]. While the two energies differ in general, they are equivalent under the MDM framework (see Sec. 3.1).
@@ 2.3 MDM as Geodesic Curve
To understand the relationship between MDM and kinetic energy from a geometric perspective, we draw on the geodesic interpretation [42]. This view emerges from a key geometric embedding: for each token, the conditional distribution $p_{t|1}=(p_{t|1}^{1},...,p_{t|1}^{D})\in\Delta^{D-1}$ can be mapped isometrically onto the unit sphere via square-root parameterization:
@@
Through this embedding, [42] proved geometrically that the per-token MDM conditional flow Eq. (3) corresponds to geodesic motion on $\mathbb{S}^{D-1}$—that is, movement along the great circle connecting the masked initial state $y_{0}$ and the target distribution $y_{1}$. The interpolation schedule $\gamma_{t}$ (a smooth increasing function satisfying $\gamma_{0}=0$, $\gamma_{1}=1$) governs the temporal progression, and its derivative $\dot{\gamma}_{t}$ represents instantaneous velocity. We retain the $\gamma_{t}$ notation from Definition 2.1, as it also serves as the weight function for energy minimization as detailed in Sec. 3. A more comprehensive and intuitive introduction to the geodesic curve and its relationship with MDM is provided in Appendix B.
This geometric framework unifies MDM sampling dynamics with energy minimization principles, since geodesics simultaneously minimize both path length and energy functionals [1]. To formalize this link, we define the weighted geodesic energy, which MDMs minimize under suitable scheduling.
@@ 3.1 Equivalence of Energies in MDMs
As discussed in Sec. 2, while $\mathcal{E}_{k}$ and $\mathcal{E}_{c}$ share a similar structure, their inherent differences between marginal and conditional formulations typically lead to divergent values. However, our proof in Appendix D.1 reveals that by decomposing the concrete score in the rate matrix of MDM (see Eq. (2)) into temporal components and clean-data conditional probabilities [44] and leveraging the inherent simple closed-form of MDM’s conditional rate matrix (characterized in Appendix C.3), equivalence between $\mathcal{E}_{k}$ and $\mathcal{E}_{c}$ are established, even in high-dimensional regimes.
@@
Our proof in Appendix D.2 reveals that this equivalence originates from the token-wise decomposition of $\mathcal{E}_{c}$ (see Appendix C.4) – a property inherently enabled by MDM’s per-token structure of $p_{t|1}$ and $Q_{t|1}$. This extends the DFM framework [37] beyond its original one-dimensional geodesic-kinetic energy correspondence. Moreover, more flexible $\gamma_{t}$ choices are considered in our settings while energy functionals in [37] use a fixed weight function $\gamma_{t}=t$. Therefore, our Theorem 3.2 establishes interpretations applicable to real-world text generation with adaptive weighting schemes.
@@ 3.2 MDMs with the Optimal Mask Schedule Are Energy Minimizers
While this relationship was geometrically established in [42], demonstrating the relationship between the interpolation schedule of the geodesic curve and the mask schedule of MDM that generates the curve, its profound implications for energy minimization remain uncharacterized – a gap our subsequent lemma addresses.
@@
Our proof in Appendix D.4 demonstrates through energy-theoretic analysis that Condition 3.4 not only guarantees generation of minimal-length geodesic paths (thereby providing a proof for existing geometric conclusions [42] in an alternative perspective) but also formally establishes the attainment of minimal geodesic energy. These results enrich our understanding of MDM optimality. One may further be curious about the practical implications: does discretizing the continuous trajectory into a finite-step sampling process affect the attainment of optimality? We address this in Appendix D.5, demonstrating that the discretized trajectory remains optimal in a well-defined sense. Therefore, our subsequent discussion will continue to focus on the theoretical continuous case.
@@ 3.3 Energy-Inspired Fast Samplers
However, direct optimization over the space of all possible schedules remains intractable due to its infinite-dimensional nature. Existing approaches therefore rely on a small set of manually designed schedules, such as linear ($\alpha_{t}=t$[8, 26, 4], sine ($\alpha_{t}=\sin\left(\frac{\pi}{2}t\right)$[36], or squared sine schedules ($\alpha_{t}=\sin^{2}\left(\frac{\pi}{2}t\right)$[17]. To bridge the gap between theoretical flexibility and practical feasibility, we propose parameterizing $\gamma_{t}$ as the cumulative distribution function (CDF) of a beta distribution with two parameters $(a,b)$:
@@ 4 Experiments
We evaluate our method using LLaDA 8B [43], an open-source MDM that achieves performance comparable to modern large language models such as LLaMA3 [31]. We select six representative tasks: MBPP [7], HumanEval [9], BBH [28], GSM8K [10], Hendrycks Math [11] and Minerva Math [18]. These benchmarks comprehensively assess the model’s capabilities in general reasoning, mathematical problem solving, and code generation. Please refer to Appendix E.2 for more experimental details.
@@
Figure 5: Performance evaluation of energy-optimized schedules on LLaDA 8B [43]. Each panel corresponds to a distinct benchmark. The x-axis displays sampling steps on a logarithmic scale, while the y-axis quantifies task performance, where higher values denote superior generation quality. Results on benchmarks where beta-parameterized schedules exhibit comparable yet not better performance are provided in Appendix. E.3.
@@ 5 Related Work
Mask Diffusion Models. MDMs [4, 8, 34, 44, 36, 38] have established themselves as a prominent class of generative models for discrete data. While substantial progress has been made in understanding MDM training dynamics through theoretically equivalent objective formulations [12, 38, 36, 44] and parameterization strategies [8, 14, 21, 20, 19], the sampling process remains relatively underexplored. Existing efforts primarily concentrate on developing advanced discrete sampling algorithms, including Tweedie $\tau$-sampling [3, 6, 26], k-Gillespie methods [48], and higher-order solvers [45], along with discretizing time step optimization techniques [35] and distillation-based acceleration [33, 46]. Notably, the critical mask schedules governing sampling trajectories have not received systematic investigation. Current implementations [8, 26, 4, 17, 36] typically employ manually designed schedules (predominantly linear), even in state-of-the-art MDM-based large language models [43]. This underscores the necessity for principled task-adaptive schedule optimization, which constitutes our primary contribution.
@@
Energy and Geodesic Perspective. The connection between kinetic energy minimization [2, 13] and efficient sampling via optimal transport trajectories [24, 22, 25, 29] has been well-established in continuous settings. However, existing literature predominantly examines continuous diffusion processes, limiting direct applicability to discrete domains. Recent geometric analyses [42] reveal intrinsic links between MDM probability paths and geodesic curves under specific interpolation schedules, motivating our exploration of geodesic perspective as a bridge between MDMs and kinetic principles. However, the energy perspective—which plays a vital role in our optimal transport framework—is absent in [42], thus preventing it from establishing the optimality of Condition 3.4 in the context of sampling efficiency. The most relevant work [37] introduces energy perspectives to Discrete Flow Matching (DFM), a distinct discrete probability flow (DPF) variant. Crucially, their framework decouples probability flow and rate matrix optimization - an approach incompatible with MDMs where $\alpha_{t}$ jointly governs both components. Despite this architectural constraint, we demonstrate through Theorem 3.6 that MDMs inherently achieve optimal rate selection by leveraging a key lemma from [37] (see Appendix D.6). Our introduced $\gamma_{t}$ interpolation schedule also provides novel theoretical insights by establishing that every mask schedule optimizes a corresponding energy functional, thereby justifying task-specific schedule tuning - a capability absent in [37] where formulations reduce to the $\gamma_{t}=t$ special case. Furthermore, the training loss invariance to mask schedules represents a unique MDM property distinguishing it from DFM and conventional DPF frameworks, enabling exclusive post-training schedule optimization within our MDM paradigm.
@@ Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. 92470118); Beijing Natural Science Foundation (No. L247030). The work was partially done at the Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education.

### v2

@@ 1 Introduction
Masked diffusion models (MDMs) [sohl2015deep, austin2021structured, sedd, radd, mdlm, shi2024simplified] have emerged as a powerful class of generative models for discrete data. By reversing a stochastic masking process, MDMs iteratively generate sequences through a series of unmasking steps, guided by learned denoising functions. This simple yet flexible architecture has shown promising empirical performance across text generation [llada], protein generation [protein1, protein2], and image generation [chang2022maskgit, chang2023muse, you2025effective].
Despite their empirical success, the underlying principles that govern the sampling efficiency of MDMs—particularly in few-step regimes—remain poorly understood. Existing works typically adopt manually designed mask schedules (e.g., linear or sine) without theoretical justification. In contrast, in continuous domains [KineticOptimal, albergo2023building, liu2023flow] and discrete flow matching [DiscretePaths], recent work has drawn deep connections between diffusion models and optimal transport, motivating questions such as: Can MDMs be similarly understood through a similar lens? Can we characterize optimal sampling schedules in a principled way? How do these schedules relate to the geometry of the underlying probability space?
@@ 2.1 Discrete Probability Flows and Masked Diffusion Models
Discrete Probability Flows (DPFs). DPFs [dieleman2022continuous, campbell2022continuous, DiscreteFlowMatching] define continuous-time transformations over structured distributions in finite state spaces. In text generation tasks, a state $z=(z^{1},...,z^{n})\in\mathcal{D}^{n}$ typically denotes a token sequence of fixed length $n$ drawn from a vocabulary $\mathcal{D}$ of size $d$. Formally, DPFs introduce a family of parameterized distributions $(p_{t}(z))_{t\in[0,1]}$ that interpolate between a tractable base distribution $p_{0}(z)$ (e.g., a uniform distribution) and a target distribution $p_{1}(z)$.
A DPF is governed by a time-dependent transition rate matrix $(Q_{t})_{t\in[0,1]}$, [^1] We adopt the MDM terminology rather than the “velocity functions” used in the flow matching literature. which specifies transition probabilities between states. However, the reverse implication does not hold: different rate matrices can induce the same DPF [DiscreteFlowMatching]. We will omit the range of the time index $t\in[0,1]$ for brevity.
Masked Diffusion Models (MDMs). MDMs [sohl2015deep, austin2021structured] represent a subclass of DPFs built upon absorbing Markov chains. These models capture two temporal processes: a masking process (operating backward in time, $t=1\rightarrow 0$) and its reverse, the unmasking process ($t=0\rightarrow 1$) as follows.
@@
$p_{t|1}(x^{i}|x_{1})=(1-\alpha_{t})\delta_{[\mathrm{M}]}(x^{i})+\alpha_{t}\delta_{x_{1}^{i}}(x^{i}),$ (3)
@@
as shown in existing work [radd, shi2024simplified]. Existing methods model the likelihood ratio $\frac{p_{t}(x)}{p_{t}(z)}$ in Eq. (2)—also known as the concrete score [meng2022concrete, lou2023discrete] or its equivalent formulations [shi2024simplified, mdlm, radd]—and optimize evidence lower bounds (ELBO) of the log-likelihood [sohl2015deep]. In particular, the ELBO is invariant to the choice of $\alpha_{t}$ [kingma2021variational, shi2024simplified, mdlm] by change of variables. We provide proofs for Eq. (4) in Appendix C.1 and the invariance of ELBO in Appendix C.2 for completeness.
@@ 2.2 Kinetic and Conditional Kinetic Energy
Kinetic energy [mccann1997convexity, villani2021topics, KineticOptimal, DiscretePaths] provides a principled framework for quantifying the transport cost of probability flows, forming an optimization objective that yield improved sampling trajectories.
@@
The function $\gamma_{t}$ is a smooth, strictly increasing schedule satisfying the boundary conditions $\gamma_{0}=0$ and $\gamma_{1}=1$. Different choices of $\gamma_{t}$ implement various temporal weighting schemes for the DPF. This generalizes the unweighted kinetic energy [DiscretePaths], where $\gamma_{t}=t$, and facilitates our theoretical analysis.
The quadratic dependence on the transition rates $Q_{t}(z,x)$ in Eq. (5) mirrors the classical velocity-squared form of kinetic energy. As $Q_{t}$ controls sampling behavior, minimizing $\mathcal{E}_{k}$ seeks minimal-cost sampling paths between distributions [FlowMatching, albergo2023building, liu2023flow], a central goal in efficient generative modeling and few-step sampling [mccann1997convexity, villani2021topics, KineticOptimal].
@@
Definition 2.2 extends Definition 2.1 to the conditional setting, in line with analogous energy formulations in the continuous domain [KineticOptimal]. While the two energies differ in general, they are equivalent under the MDM framework (see Sec. 3.1).
@@ 2.3 MDM as Geodesic Curve
To understand the relationship between MDM and kinetic energy from a geometric perspective, we draw on the geodesic interpretation [jo2025]. This view emerges from a key geometric embedding: for each token, the conditional distribution $p_{t|1}=(p_{t|1}^{1},...,p_{t|1}^{D})\in\Delta^{D-1}$ can be mapped isometrically onto the unit sphere via square-root parameterization:
@@
Through this embedding, [jo2025] proved geometrically that the per-token MDM conditional flow Eq. (3) corresponds to geodesic motion on $\mathbb{S}^{D-1}$—that is, movement along the great circle connecting the masked initial state $y_{0}$ and the target distribution $y_{1}$. The interpolation schedule $\gamma_{t}$ (a smooth increasing function satisfying $\gamma_{0}=0$, $\gamma_{1}=1$) governs the temporal progression, and its derivative $\dot{\gamma}_{t}$ represents instantaneous velocity. We retain the $\gamma_{t}$ notation from Definition 2.1, as it also serves as the weight function for energy minimization as detailed in Sec. 3. A more comprehensive and intuitive introduction to the geodesic curve and its relationship with MDM is provided in Appendix B.
This geometric framework unifies MDM sampling dynamics with energy minimization principles, since geodesics simultaneously minimize both path length and energy functionals [spivak1970comprehensive]. To formalize this link, we define the weighted geodesic energy, which MDMs minimize under suitable scheduling.
@@ 3.1 Equivalence of Energies in MDMs
As discussed in Sec. 2, while $\mathcal{E}_{k}$ and $\mathcal{E}_{c}$ share a similar structure, their inherent differences between marginal and conditional formulations typically lead to divergent values. However, our proof in Appendix D.1 reveals that by decomposing the concrete score in the rate matrix of MDM (see Eq. (2)) into temporal components and clean-data conditional probabilities [radd] and leveraging the inherent simple closed-form of MDM’s conditional rate matrix (characterized in Appendix C.3), equivalence between $\mathcal{E}_{k}$ and $\mathcal{E}_{c}$ are established, even in high-dimensional regimes.
@@
Our proof in Appendix D.2 reveals that this equivalence originates from the token-wise decomposition of $\mathcal{E}_{c}$ (see Appendix C.4) – a property inherently enabled by MDM’s per-token structure of $p_{t|1}$ and $Q_{t|1}$. This extends the DFM framework [DiscretePaths] beyond its original one-dimensional geodesic-kinetic energy correspondence. Moreover, more flexible $\gamma_{t}$ choices are considered in our settings while energy functionals in [DiscretePaths] use a fixed weight function $\gamma_{t}=t$. Therefore, our Theorem 3.2 establishes interpretations applicable to real-world text generation with adaptive weighting schemes.
@@ 3.2 MDMs with the Optimal Mask Schedule Are Energy Minimizers
While this relationship was geometrically established in [jo2025], demonstrating the relationship between the interpolation schedule of the geodesic curve and the mask schedule of MDM that generates the curve, its profound implications for energy minimization remain uncharacterized – a gap our subsequent lemma addresses.
@@
Our proof in Appendix D.4 demonstrates through energy-theoretic analysis that Condition 3.4 not only guarantees generation of minimal-length geodesic paths (thereby providing a proof for existing geometric conclusions [jo2025] in an alternative perspective) but also formally establishes the attainment of minimal geodesic energy. These results enrich our understanding of MDM optimality. One may further be curious about the practical implications: does discretizing the continuous trajectory into a finite-step sampling process affect the attainment of optimality? We address this in Appendix D.5, demonstrating that the discretized trajectory remains optimal in a well-defined sense. Therefore, our subsequent discussion will continue to focus on the theoretical continuous case.
@@ 3.3 Energy-Inspired Fast Samplers
However, direct optimization over the space of all possible schedules remains intractable due to its infinite-dimensional nature. Existing approaches therefore rely on a small set of manually designed schedules, such as linear ($\alpha_{t}=t$[austin2021structured, lou2023discrete, sohl2015deep], sine ($\alpha_{t}=\sin\left(\frac{\pi}{2}t\right)$[mdlm], or squared sine schedules ($\alpha_{t}=\sin^{2}\left(\frac{\pi}{2}t\right)$[han2022ssd]. To bridge the gap between theoretical flexibility and practical feasibility, we propose parameterizing $\gamma_{t}$ as the cumulative distribution function (CDF) of a beta distribution with two parameters $(a,b)$:
@@ 4 Experiments
We evaluate our method using LLaDA 8B [llada], an open-source MDM that achieves performance comparable to modern large language models such as LLaMA3 [dubey2024llama]. We select six representative tasks: MBPP [metric:mbpp], HumanEval [metric:humaneval], BBH [metric:bbh], GSM8K [metric:gsm8k], Hendrycks Math [metric:math] and Minerva Math [lewkowycz2022solving]. These benchmarks comprehensively assess the model’s capabilities in general reasoning, mathematical problem solving, and code generation. Please refer to Appendix E.2 for more experimental details.
@@
Figure 5: Performance evaluation of energy-optimized schedules on LLaDA 8B [llada]. Each panel corresponds to a distinct benchmark. The x-axis displays sampling steps on a logarithmic scale, while the y-axis quantifies task performance, where higher values denote superior generation quality. Results on benchmarks where beta-parameterized schedules exhibit comparable yet not better performance are provided in Appendix. E.3.
@@ 5 Related Work
Mask Diffusion Models. MDMs [sohl2015deep, austin2021structured, sedd, radd, mdlm, shi2024simplified] have established themselves as a prominent class of generative models for discrete data. While substantial progress has been made in understanding MDM training dynamics through theoretically equivalent objective formulations [kingma2021variational, shi2024simplified, mdlm, radd] and parameterization strategies [austin2021structured, campbell2022continuous, xie2022vector, sun2022score, meng2022concrete], the sampling process remains relatively underexplored. Existing efforts primarily concentrate on developing advanced discrete sampling algorithms, including Tweedie $\tau$-sampling [gillespie2001approximate, eweinan, lou2023discrete], k-Gillespie methods [zhao2025informed], and higher-order solvers [FastSolver], along with discretizing time step optimization techniques [JumpYourSteps] and distillation-based acceleration [kou2024cllms, xu2025show]. Notably, the critical mask schedules governing sampling trajectories have not received systematic investigation. Current implementations [austin2021structured, lou2023discrete, sohl2015deep, han2022ssd, mdlm] typically employ manually designed schedules (predominantly linear), even in state-of-the-art MDM-based large language models [llada]. This underscores the necessity for principled task-adaptive schedule optimization, which constitutes our primary contribution.
@@
Energy and Geodesic Perspective. The connection between kinetic energy minimization [mccann1997convexity, villani2021topics] and efficient sampling via optimal transport trajectories [FlowMatching, albergo2023building, liu2023flow, FlowThroughTransport] has been well-established in continuous settings. However, existing literature predominantly examines continuous diffusion processes, limiting direct applicability to discrete domains. Recent geometric analyses [jo2025] reveal intrinsic links between MDM probability paths and geodesic curves under specific interpolation schedules, motivating our exploration of geodesic perspective as a bridge between MDMs and kinetic principles. However, the energy perspective—which plays a vital role in our optimal transport framework—is absent in [jo2025], thus preventing it from establishing the optimality of Condition 3.4 in the context of sampling efficiency. The most relevant work [DiscretePaths] introduces energy perspectives to Discrete Flow Matching (DFM), a distinct discrete probability flow (DPF) variant. Crucially, their framework decouples probability flow and rate matrix optimization - an approach incompatible with MDMs where $\alpha_{t}$ jointly governs both components. Despite this architectural constraint, we demonstrate through Theorem 3.6 that MDMs inherently achieve optimal rate selection by leveraging a key lemma from [DiscretePaths] (see Appendix D.6). Our introduced $\gamma_{t}$ interpolation schedule also provides novel theoretical insights by establishing that every mask schedule optimizes a corresponding energy functional, thereby justifying task-specific schedule tuning - a capability absent in [DiscretePaths] where formulations reduce to the $\gamma_{t}=t$ special case. Furthermore, the training loss invariance to mask schedules represents a unique MDM property distinguishing it from DFM and conventional DPF frameworks, enabling exclusive post-training schedule optimization within our MDM paradigm.
@@ Acknowledgement
This work was supported by the National Natural Science Foundation of China (No. 92470118); the Beijing Natural Science Foundation (No. L247030); the Beijing Nova Program (No. 20230484416); the Public Computing Cloud of Renmin University of China; the fund for building world-class universities (disciplines) of Renmin University of China; and the Huawei Research Fund.